Selena Company has two products: A and B. The company uses activity-based costing. The estimated total cost and expected activity for each of the company's three activity cost pools are as follows:

Expected Activity
Activity Cost Pool Estimated Cost Product A Product B Total
Activity 1 $17,600 800 300 1,100
Activity 2 12,000 500 200 700
Activity 3 26,000 800 400 1,200



The activity rate under the activity-based costing system for Activity 3 is closest to which of the following?
Answer #1

Activity rate = Estimated overhead/Total expected activity

= 26,000/1200

= 21.67

Comment if you face any issues.

Thank you
Answer #2

o find the activity rate for Activity 3 under the activity-based costing system, we need to divide the estimated cost for Activity 3 by the total activity for all products combined. The formula for calculating the activity rate is as follows:

Activity Rate = Estimated Cost / Total Activity

Let's calculate the activity rate for Activity 3:

Estimated Cost for Activity 3 = $26,000 Total Activity for Activity 3 (Product A + Product B) = 800 + 400 = 1,200

Activity Rate for Activity 3 = $26,000 / 1,200 = $21.67 (approximately)

So, the activity rate under the activity-based costing system for Activity 3 is closest to $21.67.


answered by: mervetokaz